OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). A. y = -4x + 1 B. y = -1/2 x + 9 C. y = 4x + 1 D. y = -2x + 9

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

1st... find slope (-1-1) / (5-4) = -2/1 = -2 So the slope is -2. Next step... find y-intercept. This is where the line crosses the y-axis. To find it... use this equation y=mx+b m= slope = -2 b= y-intercept Using slope and one of the coordinates you are given, find b I'll use the (5, -1) one y=mx+b -1=-2(5)+b b= 9 So slope is -2, b is 9. That's all you need! Plug in those in the equation: y=mx+b y= -2x+9